# Additive Sink Subtraction

### v2

@@ Abstract
Subtraction games are a classical topic in Combinatorial Game Theory. An early result of Golomb (1966) shows that every subtraction game with a finite move set has an eventually periodic nim-sequence, but the known proof yields only an exponential upper bound on the period length. Even three-move games are not yet fully understood. Nevertheless, Flammenkamp (1997) conjectures a striking classification: non-additive rulesets exhibit linear period lengths of the form “the sum of two moves”, where the choice of which two moves displays fractal-like behavior, while additive sets $S=\{a,b,a+b\}$ have purely periodic outcomes with linear or quadratic period lengths. Despite early attention in Winning Ways (1982), the general additive case remains open. We analyze a dual winning convention, called sink subtraction. Unlike the standard wall convention, where moves to negative positions are forbidden, the sink convention declares a player the winner upon moving to a non-positive position. We show that additive sink subtraction admits a complete solution: its nim-sequence is purely periodic with an explicit linear or quadratic period formula. We further conjecture a nimber duality between additive sink subtraction and classical wall subtraction.
@@ 1 Introduction
For example, if $S=\{2,5\}$, then starting from $x=0$, ‘wall’ has a purely periodic nim-sequence of period length 7,
@@
while ‘sink’, starting from $x=1$, begins
@@
$S(m,\delta)=\{m,m+\delta,2m+\delta\},$ (2)
where $m,\delta\in\mathbb{N}$. There is a conjecture from Flammenkamp’s PhD thesis [3] concerning the periodicity of additive wall subtraction, and this partially paraphrases the treatment in Winning Ways [2], where the linear-period-length part of the conjecture has been proved. As far as the authors know, the non-linear (quadratic) part has not yet been established. [^1] Non-additive three-move games are conjectured to have period lengths of the form “the sum of two moves”, where “which two moves?” appear to satisfy some fractal patterns [3]. Let us state our main result.
@@
$v_{\rm wall}\bigl([x]_{p}\bigr)=\sigma\bigl(v_{\rm sink}\bigl([s_{2}-x]_{p}\bigr)\bigr),$ (3)
@@ 2 A proof of the main result
It remains to justify the general case of item (ii), for a fixed $k$, and when $\delta=2nm+k$, for $n\geq 1$. For each increase in $n$, the period length increases by $4m^{2}/{\gcd(m,\delta)}$, and this is due to the fact that each block is increased by exactly $4m$ positions via the two concatenated factors $0^{m}1^{m}$ and $2^{m}3^{m}$; see Figure 4. The justification is similar to case (i), but here it invokes another quadratic term. Apart from this observation, the mex verification remains the same as for the case $n=1$. This concludes the proof of Theorem 3.
@@ 3 Final remarks
On a more philosophical matter, with respect to the simplicity of the proof of Theorem 2, one might ponder the value of understanding the fine details of a small sub-family of subtraction games. However this type of result has been on the list of conjectures for a long time, and our conjectured reciprocity with the wall convention requires further research. For example, Richard Guy writes in the first problem, in the first edition of Unsolved problems in combinatorial games [5] “Subtraction games are known to be periodic. Investigate the relationship between the subtraction set and the length and structure of the period.”; see also [7]. The second part of the conjecture for three move games [3, 8] concerns the non-additive three-move rulesets. Although they all are conjectured to have period length “a sum of two moves”, and any such instance would individually be easily verified, the full classification appears to satisfy some fractal structure [3], and has not yet been fully understood. See also the recent survey on subtraction games [6].

### v3

@@ Abstract
Subtraction games are a classical topic in Combinatorial Game Theory. An early result of Golomb (1966) shows that every subtraction game with a finite move set has an eventually periodic nim-sequence, but the known proof yields only an exponential upper bound on the period length. Even three-move games are not yet fully understood. Nevertheless, Flammenkamp (1997) conjectures a striking classification: non-additive rulesets have linear period lengths of the form “the sum of two moves”, where the choice of which two moves displays fractal-like behavior, while additive sets $S=\{a,b,a+b\}$ have purely periodic outcomes with linear or quadratic period lengths. Despite early attention in Winning Ways (1982), the general additive case remains open. We analyze a dual winning convention, called sink subtraction. Unlike the standard wall convention, where moves to negative positions are forbidden, the sink convention declares a player the winner upon moving to a non-positive position. We show that additive sink subtraction admits a complete solution: its nim-sequence is purely periodic with an explicit linear or quadratic period formula. We further conjecture a nimber duality between additive sink subtraction and classical wall subtraction.
@@ 1 Introduction
For example, if $S=\{2,5\}$, then starting from $x=0$, wall has a purely periodic nim-sequence of period length 7,
@@
while sink, starting from $x=1$, begins
@@
$S=\{s_{1},s_{2},s_{1}+s_{2}\},$ (2)
where $s_{1},s_{2}\in\mathbb{N}$, $s_{1}<s_{2}$. There is a conjecture from Flammenkamp’s PhD thesis [3] concerning the periodicity of additive wall subtraction, and this partially paraphrases the treatment in Winning Ways [2], where the linear-period-length part of the conjecture has been explained (without a proof). As far as the authors know, the non-linear (quadratic) part has not yet been fully solved in the literature, but conjectured also in [9]. [^1] Non-additive three-move games are conjectured to have period lengths of the form “the sum of two moves”, where “which two moves?” appear to satisfy some fractal patterns [3]. Recently a subset of the authors completely determine the closed formulas of the nimbers in the primitive quadratic regime [6] (whenever $a<\delta<2a$ and $\gcd{a,\delta}=1$), by applying elementary combinatorial number theory to the $\mathcal{P}$-position bracket expression from [2] (see also Section 3).
Let us state our main result. In the statement we use a different representation of (2).
@@
$v_{\rm wall}\bigl([x]_{p}\bigr)=\sigma\bigl(v_{\rm sink}\bigl([-s_{2}-x]_{p}\bigr)\bigr),$ (3)
@@
Let us remark that the well-known Ferguson pairing principle [2] fails to hold in sink subtraction.
Remark 6.
Let us first reiterate the classical induction proof that, in the wall convention, $v(x)=0$ if and only if $v(x+m)=1$, where $m=\min S$. We then explain how this can fail in the sink convention.
Suppose $v(x)=0$ but $v(x+m)>1$. Then there exists a move $s$ such that v(x+ms)=1v(x+m-s)=1. Since $v(x)=0$, we have v(xs)>0v(x-s)>0 for all wall-feasible $s$. Hence subtracting $m$ from a heap of size x+msx+m-s does not yield a 0-position, which contradicts the induction hypothesis that every value “1” on a smaller heap must be $m$-preceded by a “0”.
Conversely, suppose $v(x+m)=1$ but $v(x)>0$, which implies $v(x)>1$. Then there exists $s$ such that v(xs)=0v(x-s)=0. By induction, it follows that v(x+ms)=1v(x+m-s)=1, which contradicts the assumption that $v(x+m)=1$.
In the sink convention, the base case fails; see for example Example 5, due to the presence of additional terminal zeroes. Observe that $v(5)=1$, but $v(3)\neq 0$. Indeed, the step “By induction, then v(x+ms)=1v(x+m-s)=1” does not apply, since in this case x+ms=0x+m-s=0, which belongs to the sink. Similarly, $v(8)=0$ but $v(10)\neq 1$. Here $v(8)$ is the first non-trivial zero-value, yet $v(10)\neq 1$ because $v(5)=1$. These violations of Ferguson’s principle propagate to larger heaps. Thus, in the sink convention, one can no longer deduce the $1$-values from the 0-values, or vice versa.
Note that in the case of sink, the above conjecture, together with Ferguson’s pairing principle, instead implies a pairing of “1”s and “2”s: $v_{\mathrm{sink}}(x)=1$ if and only if $v_{\mathrm{sink}}(x-m)=2$.
@@ 2 A proof of the main result
It remains to justify the general case of item (ii), for a fixed $k$, and when $\delta=2nm+k$, for $n\geq 1$. For each increase in $n$, the period length increases by $4m^{2}/{\gcd(m,\delta)}$, and this is due to the fact that each block is increased by exactly $4m$ positions via the two concatenated factors $1^{m}2^{m}$ and $3^{m}0^{m}$. Precisely, for each increase in $n$, insert $1^{m}2^{m}$ just after each $A$-prefix and insert $3^{m}0^{m}$ exactly between the last zero-factor and the last three-factor (independently of whether a $B$- or $C$-block); see Figure 4. The mex-justification is similar to case (i), but here it invokes another quadratic term. Apart from this observation, the mex verification remains the same as for the case $n=1$. This concludes the proof of Theorem 3.
@@ 3 Final remarks
On a more philosophical matter, with respect to the simplicity of the proof of Theorem 2, one might ponder the value of understanding the fine details of a small sub-family of subtraction games. However this type of result has been on the list of conjectures for a long time, and our conjectured reciprocity with the wall convention requires further research. For example, Richard Guy writes in the first problem, in the first edition of Unsolved problems in combinatorial games [5] “Subtraction games are known to be periodic. Investigate the relationship between the subtraction set and the length and structure of the period.”; see also [8]. The second part of the conjecture for three move games [3, 9] concerns the non-additive three-move rulesets. Although they all are conjectured to have period length “a sum of two moves”, and any such instance would individually be easily verified, the full classification appears to satisfy some fractal structure [3], which has not yet been fully understood. See also the recent survey on subtraction games [7].
There is a closed formula expression for the $\mathcal{P}$-positions of additive wall subtraction in [2, 3]; and hence, by (1), it applies also to those of additive sink subtraction. Namely the set of wall $\mathcal{P}$-positions is generated by, for all $n\geq 0$,
$n+\left\lfloor\frac{n}{m}\right\rfloor m+\left\lfloor\frac{2n}{2m+\delta-d}\right\rfloor(m+\delta)\ \text{ or }\ n+\left\lfloor\frac{n}{m}\right\rfloor m+\left\lfloor\frac{2n}{\delta+d}\right\rfloor(m+\delta),$
depending on whether $d\leq m$ or not. To emphasize that the first (second) expression gives a linear (quadratic) period length, we rewrite this as
$w_{n}=n+\left\lfloor\frac{n}{m}\right\rfloor m+\left\lfloor\frac{2n}{km}\right\rfloor(m+\delta)\ \text{ or }\ w_{n}=n+\left\lfloor\frac{n}{m}\right\rfloor m+\left\lfloor\frac{2n}{\ell m+2d}\right\rfloor(m+\delta),$
with $k=(2m+\delta-d)/m\in\mathbb{N}$ and $\ell=(\delta-d)/m\in\mathbb{N}$, respectively. To illustrate, we get the following sequences for the parametrizations $m=2,\delta=d=1$ and $m=2,\delta=d=3$, respectively.
$n$ 0 1 2 3 4 5 6 7 8 9 10 11
$w_{n}(2,1)$ 0 1 7 8 14 15 21 22 28 29 35 36
$s_{n}(2,1)$ 5 6 12 13 19 20 26 27 33 34 40 41
$n$ 0 1 2 3 4 5 6 7 8 9 10 11
$w_{n}(2,3)$ 0 1 4 10 13 14 22 23 26 32 35 36
$s_{n}(2,3)$ 7 8 11 17 20 21 29 30 33 39 42 43
Here, we have included, as a second data row, the corresponding sink-shifted $\mathcal{P}$-positions.
Let us use the conjecture in Equation 3 to directly compute all the nimbers of both sink and wall for the first ruleset, $S=\{2,3,5\}$ using also Ferguson’s pairing principle. By the conjecture, from the $\mathcal{P}$-positions, we derive the nimber two positions for both conventions.
For all $x\leq 12$, if $v_{w}(x)=0$, then $v_{s}(14-3-x)=2$. Hence we get the sink $2$-sequence: $3,4,10,11,\ldots$
Similarly, for all $x\leq 17$, if $v_{s}(x)=0$, then $v_{w}(14+3-x)=2$. Hence we get the wall $2$-sequence: $4,5,11,12,\ldots$
Next, the nimber one positions in the wall convention are obtained via the classic Ferguson’s pairing principle as $2,3,9,10,\ldots$. Now, the nimber three positions in the wall convention must be the remaining positions, namely: $6,13,20,\ldots$
Similarly in the sink convention, the conjecture gives the postions of the nimber ones by the formula 143x14-3-x, where position $v_{w}(x)=1$. Thus, the sequence starts: $1,2,8,9,\ldots$ And the nimber three positions in the sink convention must be the remaining positions, namely: $5,12,19,\ldots$
Note that all these values are obtained without any mex computation, from the classic $\mathcal{P}$-position formula in the wall convention. Similarly one can fill in the periodic nimber positions for the ruleset $\{2,5,7\}$, without using any mex-rule.
Let us summarize these findings. Let $S=\{s_{1},s_{2},s_{3}\}$ be an additive ruleset, with $s_{1}<s_{2}<s3$.
  1. 1.
    By wall-sink outcome translation, $v_{w}(x)=0$ if and only if $v_{s}(x+s_{3}+1)=0$.
  2. 2.
    By Ferguson, wall zeros and ones are paired by the smallest move: $v_{w}(x)=0$ if and only $v_{w}(x+s_{1})=1$.
  3. 3.
    By our conjecture, given a wall value at position $x$, then $\sigma(v_{w}(x))=v_{s}(-s_{2}-x)$.
  4. 4.
    By our conjecture, given a sink value at position $x$, then $\sigma(v_{s}(x))=v_{w}(s_{2}-x)$.
  5. 5.
    Altogether by using the known formula of the $\mathcal{P}$-positions for the wall convention, we can therefore compute the positions of all nimbers in $\{0,1,2\}$ for both conventions.
  6. 6.
    Since the ruleset has exactly three moves, we may conclude that, for each convention, the remaining positions have value three, respectively.
From our duality conjecture together with Ferguson’s pairing principle, let us derive explicit descriptions of all nimber classes in both conventions in terms of the wall $\mathcal{P}$-position sequence $(w_{n})$ and the period $p$ of the outcome patterns.
For the wall convention we get the sets of positions with nimbers smaller than three as:
$\mathcal{W}_{0}$ $=\{\,w_{n}:n\geq 0\,\},$
$\mathcal{W}_{1}$ $=\{\,w_{n}+s_{1}:n\geq 0\,\},$
$\mathcal{W}_{2}$ $=\{\,[\,-w_{n}-s_{1}\,]_{p}+kp:n\geq 0,\ k\geq 0\,\}.$
And similarly for the sink convention:
$\mathcal{S}_{0}$ $=\{\,w_{n}+s_{3}+1:n\geq 0\,\},$
$\mathcal{S}_{1}$ $=\{\,[\,-s_{3}-w_{n}\,]_{p}+kp:n\geq 0,\ k\geq 0\,\},$
$\mathcal{S}_{2}$ $=\{\,[\,-s_{2}-w_{n}\,]_{p}+kp:n\geq 0,\ k\geq 0\,\}.$
As before, $x]_{p}$ denotes the smallest nonnegative representative of $x$ modulo $p$. In either case the nimber three positions are the remaining ones. Given our conjecture, in the case of “primitive quadratic” the nimber classes are also obtained from the wall bracket formulas [6] by linear transformations and reduction modulo $p$.